# Are We Done with MMLU?

### v2

@@ Abstract
Maybe not. We identify and analyse errors in the popular Massive Multitask Language Understanding (MMLU) benchmark. Even though MMLU is widely adopted, our analysis demonstrates numerous ground truth errors that obscure the true capabilities of LLMs. For example, we find that 57% of the analysed questions in the Virology subset contain errors. To address this issue, we introduce a comprehensive framework for identifying dataset errors using a novel error taxonomy. Then, we create MMLU-Redux, which is a subset of 3,000 manually re-annotated questions across 30 MMLU subjects. Using MMLU-Redux, we demonstrate significant discrepancies with the model performance metrics that were originally reported. Our results strongly advocate for revising MMLU’s error-ridden questions to enhance its future utility and reliability as a benchmark. Therefore, we open up MMLU-Redux for additional annotation https://huggingface.co/datasets/edinburgh-dawg/mmlu-redux.
@@ 1 Introduction
The advent of transformer-based Large Language Models (LLMs) [1, 2, 3, 4, 5, 6, 7, 8] marked a significant advancement in generative models, enabling interaction with computing devices through natural language. This advancement rendered many earlier benchmarks and leaderboards obsolete [9, 10], leading to the compilation of more challenging and comprehensive tests. Among these benchmarks, Massive Multitask Language Understanding (MMLU) [11] has gained significant popularity: It assesses both the breadth and depth of language understanding capabilities of current LLMs across a diverse range of subjects, including mathematics, history, computer science, logic, law, etc.
However, the reliability of benchmarking results is only as robust as the quality of the dataset used. We find that, despite its popularity, MMLU suffers from numerous errors that can mislead evaluation and model comparison [12, 13]. These errors, which range from simple parsing and scraping mistakes to more complex issues related to context, interpretation, and dataset quality, compromise the reliability of MMLU as a benchmark. For example, we find that 57% of the analysed instances in the Virology subset contain errors, including the suggestion to send the American army to West Africa to prevent outbreaks of Ebola (see Fig. 1).
Therefore, in this study, we manually analyse the MMLU dataset using a novel error taxonomy to construct MMLU-Redux: 14 human experts manually assessed and re-annotated 3,000 questions across 30 subsets of MMLU. After our manual re-annotation effort, we study how the errors in MMLU impact LLM evaluation. First, we re-evaluate leading LLMs on MMLU-Redux, and found the performance metrics notably altered, changing their ranking. Furthermore, we both quantitatively and qualitatively analysed the errors to help understand how these errors impact LLM evaluation.
While MMLU-Redux provides a significant stepping stone to correcting up MMLU, reviewing the remaining 12,908 questions in MMLU is still a significant effort, highlighting the need for an automated approach. MMLU-Redux can also be leveraged as a strong benchmark for automatic error detection in NLP datasets, which would help scale up the review of benchmark datasets. Therefore, we study whether LLMs can help with error detection, using prompting techniques (i.e., In-Context Learning [14], Chain of Thoughts (CoT) [15]), Retrieval Augmented Generation (RAG) [16], and fine-tuning.We believe MMLU-Redux underscores the need for closely studying and reassessing the benchmarks used for evaluating NLP models. MMLU-Redux is available at https://huggingface.co/datasets/edinburgh-dawg/mmlu-redux.
@@ 2 What is wrong with MMLU?
Figure 1: Three examples of erroneous instances from MMLU Virology, College Chemistry, and Human Aging.
The MMLU dataset has become a popular choice for evaluating the performance of NLP systems owing to its extensive coverage of many subjects collected from freely available online sources with the help of graduate and undergraduate students [11]. Despite the manual effort, MMLU still contains errors which are difficult to trace due to its under-documented annotation procedure. These errors have not yet been systematically studied, even though they were recently highlighted [12, 13].
We identify numerous errors in the MMLU questions, ranging from simple parsing mistakes – where the ground truth label differs from the correct answer in the original source (e.g., the answer from the source is B, but it is annotated as C in MMLU) – to more complex issues such as missing context. These errors appear randomly, and due to the lack of comprehensive documentation, it is difficult to trace back and identify the root causes of these problems.
This randomness and lack of traceability highlight the need for a standardised categorisation of errors to improve the reliability and accuracy of the MMLU dataset. By systematically identifying and categorising these errors, we can enhance the dataset’s quality and provide more robust benchmarks for evaluating NLP systems. Our approach involved developing a hierarchical taxonomy of errors, which we used to develop MMLU-Redux: A manual annotation of 30 subsets of MMLU, each containing 100 randomly selected samples (Section 3.1).
2.1 Error Detection Taxonomy
Figure 2: The error taxonomy used for annotating instances in MMLU — annotators first assess whether the instance is unambiguous (Question Assessment, then whether there is a single valid answer, and then whether it matches with the ground truth answer from the dataset.
We develop a hierarchical taxonomy to classify the various errors identified in MMLU into specific error types. Figure 2 illustrates our taxonomy for categorising MMLU errors, while Figure 4 provides examples of each error category. We categorise errors into two primary groups: samples with errors in the clarity of the questions (Type 1, Question Assessment) and samples with errors in the ground truth answer (Type 2, Ground Truth Verification).
@@ 2.1 Error Detection Taxonomy
(2a) No Correct Answer: None of the options correctly answer the question. This error can, for example, arise when the ground-truth options are omitted to reduce the number of options from five to four.
@@
For creating MMLU-Redux, we followed the proposed taxonomy sequentially to annotate the question. We aim to ensure comprehensive coverage of the different error types for further experiments.
2.2 Heterogeneity of Errors in MMLU Subsets
During our annotation process using the taxonomy, we observe that the types of errors found can vary substantially across subsets. Some subsets predominantly suffer from ambiguous questions, while others are mainly impacted by incorrect ground truth labels. These variations could have important implications for interpreting MMLU results in specific areas and for our ability to address these issues effectively. For instance, we have identified notable irregularities in certain subsets (a complete list of the reviewed subsets and their corresponding errors can be found in Appendix D):
@@ 2.2 Heterogeneity of Errors in MMLU Subsets
– Incorrect ground truths labels are particularly prevalent within the Virology subset. Many of the incorrect labels are for relatively simple questions, such as identifying the description of a pandemic, this suggests errors may stem from problems parsing the original datasets (in most cases the Human Virology textbook’s student resources).
@@
– The questions were sourced from textbooks and standardised college-level exams. We identified erroneous questions resulting from simple parsing errors and unknown causes. For example, questions spanning multiple lines in the original source were often parsed incorrectly, leading to a part of the question being presented as the first MMLU choice (Option A) and the exclusion of Option D. Furthermore, there were questions with ground truth labels that did not match the answers provided in the source, with no apparent cause for the discrepancy.
@@
– A major issue is the lack of specificity regarding jurisdictions. The benchmark does not clearly distinguish between different jurisdictions despite focusing on U.S. law.
@@
The above irregularities showcase some of the error patterns present in MMLU. We want to highlight one type of error that is especially challenging to catch, namely unspecified context that is needed to properly answer the question. For instance, the Professional Law dataset introduces this bias introduced by assuming the questions relate to US jurisdiction. This pattern also appears in the Professional Accounting dataset, which assumes US accounting practice. These questions may also become outdated if the law or practice changes. In general, we find several subjects to be US- and Western centric: e.g., the Virology dataset contains a question about “the Latino population”, implicitly referring to the Latino community in the US, and the Human Aging dataset discussing an unspecified survey of older adults.
Figure 3: Statistics of the error types detected in 100 randomly sampled and manually annotated data from 30 MMLU subjects — in the Virology subject, we found that 57% of the analysed instances contain some forms of errors, such as wrong ground-truth (33%), multiple correct answers (4%), or unclear questions (14%).
Figure 4: An example for every leaf node in the error taxonomy shown in Fig. 2 from the MMLU dataset.
@@ 3 MMLU-Redux: A correct MMLU Subset
We create MMLU-Redux by manually labelling a subset of MMLU questions with their corresponding error types. To this end, we follow the taxonomy introduced in Section 2.1. For more accurate annotations, we confirmed the error detection by finding the samples’ original source wherever it was available. However, at present, the correct answers suggested by the annotators are not used to replace the existing MMLU labels.
In the following, we analyse the error statistics of MMLU-Redux and use MMLU-Redux to re-evaluate the performance of LLMs. Furthermore, in Section 4, we explore the possibility of using MMLU-Redux to improve the overall quality of MMLU by automatically fixing the identified errors.
@@ 3.1 Analysis of MMLU-Redux
We present the percentage of error types in Fig. 3, with detailed numbers available in Appendix A. In our analysis, we find that more than 9% of the examples are incorrect, suggesting a substantial presence of errors in the MMLU. Especially, we find that more than 57% examples in Virology contain errors, where 30% examples have a wrong ground truth, and 15% are unclear questions. Moreover, we also observe significant error percentages in other disciplines: more than 20% examples in Logical Fallacies and College Chemistry are wrong and more than 10% examples in Professional Law, Business Ethics, Formal Logic, Human Aging, Global Facts, Machine Learning, Miscellaneous and Public Relations are wrong. Such error proportions could lead to inaccurate comparisons and invalid rankings of LLM models.
Figure 5: Exact Match (EM) for each model on all subjects combined (top left), and the seven subjects with the most errors in our MMLU-Redux data. Original EM represents performance measured across all instances regardless of error presence, while OK and Erroneous EM reflect results on correct instances and error instances, respectively.
To improve our understanding of how these errors impact the performance of models on the MMLU, we compare the performance between erroneous instances and correct instances across the seven subjects identified as having the most errors (Virology, Logical Fallacies, College Chemistry, Professional Law, Business Ethics, Formal Logic, and Human Aging) in Fig. 5.
Although the general trend indicates a performance decline among erroneous instances, we also observed cases where performance was similar or even higher in erroneous instances (Professional Law and Formal Logic). Considering that erroneous instances should intrinsically be unable to yield correct answers, this may serve as evidence of memorisation, suggesting that these MMLU instances were learned during the models’ pretraining processes.  [^2] For more detailed results including all the models and subjects, refer to https://huggingface.co/spaces/edinburgh-dawg/MMLU-Redux-EDA
@@ 3.2 Re-Evaluating the State-of-the-Art LLMs
To assess how the corrected dataset impacts the performance of existing state-of-the-art LLMs, we re-evaluate them on the five subjects with the highest number of errors (Business Ethics, Virology, Professional Law, College Chemistry, and Logical Fallacies) in MMLU-Redux.
@@
In Table 1, we compared the performance of models when using all instances of MMLU-Redux to the performance when using only correct instances without errors to see if there are any changes in the rankings due to this. The results clearly demonstrate that, at least for subjects with a high number of detected errors, these issues are significant enough to affect the results, leading to changes in model rankings. For example, in the Virology subset, Palmyra X v3 ranked 4th in overall performance when considering all instances and ranked 1st when only correct instances were used. This indicates that errors in MMLU are a critical issue, considering that MMLU is used as an important benchmark for evaluating model performance.
@@
Table 1: Comparison of model performance and ranking changes when using overall instances versus correct instances of MMLU-Redux on five most erroneous subjects.
@@
Model Business Ethics Virology Professional Law College Chemistry Logical Fallacies
Claude 3 Opus 0.86 (1) $\rightarrow$ 0.95 (2) 0.54 (9) $\rightarrow$ 0.88 (6) 0.69 (4) $\rightarrow$ 0.72 (3) 0.60 (2) $\rightarrow$ 0.72 (1) 0.90 (1) $\rightarrow$ 0.96 (3)
GPT-4o 0.85 (2) $\rightarrow$ 0.96 (1) 0.56 (1) $\rightarrow$ 0.91 (3) 0.70 (3) $\rightarrow$ 0.70 (4) 0.61 (1) $\rightarrow$ 0.71 (2) 0.89 (5) $\rightarrow$ 0.99 (1)
Gemini 1.5 Pro 0.80 (7) $\rightarrow$ 0.88 (9) 0.55 (7) $\rightarrow$ 0.84 (10) 0.67 (7) $\rightarrow$ 0.70 (5) 0.58 (6) $\rightarrow$ 0.69 (6) 0.88 (7) $\rightarrow$ 0.96 (5)
GPT-4 (0613) 0.79 (8) $\rightarrow$ 0.93 (3) 0.56 (2) $\rightarrow$ 0.88 (7) 0.71 (1) $\rightarrow$ 0.74 (1) 0.55 (9) $\rightarrow$ 0.68 (7) 0.89 (6) $\rightarrow$ 0.96 (4)
Llama 3 70b 0.83 (3) $\rightarrow$ 0.93 (4) 0.55 (8) $\rightarrow$ 0.91 (4) 0.55 (10) $\rightarrow$ 0.56 (10) 0.56 (8) $\rightarrow$ 0.67 (8) 0.85 (10) $\rightarrow$ 0.96 (6)
Gemini 1.5 Flash 0.82 (6) $\rightarrow$ 0.91 (8) 0.53 (10) $\rightarrow$ 0.88 (9) 0.58 (9) $\rightarrow$ 0.57 (9) 0.60 (3) $\rightarrow$ 0.71 (5) 0.86 (9) $\rightarrow$ 0.92 (10)
Palmyra X v3 0.83 (4) $\rightarrow$ 0.91 (6) 0.56 (4) $\rightarrow$ 0.93 (1) 0.68 (5) $\rightarrow$ 0.65 (7) 0.59 (4) $\rightarrow$ 0.71 (3) 0.90 (2) $\rightarrow$ 0.95 (7)
PaLM 2 0.83 (5) $\rightarrow$ 0.91 (7) 0.56 (5) $\rightarrow$ 0.93 (2) 0.68 (6) $\rightarrow$ 0.65 (8) 0.59 (5) $\rightarrow$ 0.71 (4) 0.90 (3) $\rightarrow$ 0.95 (8)
GPT-4 Turbo (1106) 0.78 (9) $\rightarrow$ 0.91 (5) 0.56 (3) $\rightarrow$ 0.88 (8) 0.71 (2) $\rightarrow$ 0.73 (2) 0.47 (10) $\rightarrow$ 0.60 (10) 0.86 (8) $\rightarrow$ 0.93 (9)
Mixtral 8x22b 0.74 (0) $\rightarrow$ 0.84 (10) 0.56 (6) $\rightarrow$ 0.91 (5) 0.59 (8) $\rightarrow$ 0.66 (6) 0.57 (7) $\rightarrow$ 0.64 (9) 0.90 (4) $\rightarrow$ 0.99 (2)

### v3

@@ Abstract
Maybe not. We identify and analyse errors in the popular Massive Multitask Language Understanding (MMLU) benchmark. Even though MMLU is widely adopted, our analysis demonstrates numerous ground truth errors that obscure the true capabilities of LLMs. For example, we find that 57% of the analysed questions in the Virology subset contain errors. To address this issue, we introduce a comprehensive framework for identifying dataset errors using a novel error annotation protocol. Then, we create MMLU-Redux, which is a subset of 5,700 manually re-annotated questions across all 57 MMLU subjects. We estimate that 6.49% of MMLU questions contain errors. Using MMLU-Redux, we demonstrate significant discrepancies with the model performance metrics that were originally reported. Our results strongly advocate for revising MMLU’s error-ridden questions to enhance its future utility and reliability as a benchmark. https://huggingface.co/datasets/edinburgh-dawg/mmlu-redux-2.0.
Are We Done with MMLU?
Aryo Pradipta Gema1   Joshua Ong Jun Leang1   Giwon Hong1   Alessio Devoto2 Alberto Carlo Maria Mancino2,3Rohit Saxena1Xuanli He4Yu Zhao1Xiaotang Du1 Mohammad Reza Ghasemi Madani5Claire Barale1Robert McHardy6Joshua Harris7 Jean Kaddour4Emile van Krieken1Pasquale Minervini1,8 1University of Edinburgh  2Sapienza University of Rome  3Polytechnic University of Bari 4University College London  5University of Trento  6AssemblyAI 7UK Health Security Agency  8Miniml.AI {first.last, jong2, p.minervini}@ed.ac.uk alessio.devoto@uniroma1.italberto.mancino@poliba.it mr.ghasemimadani@unitn.itjoshua.harris@ukhsa.gov.uk {xuanli.he, jean.kaddour.20, robert.mchardy.20}@ucl.ac.uk
@@ 1 Introduction
The advent of transformer-based Large Language Models (LLMs) (OpenAI, 2023; Anil et al., 2023a; Anthropic, 2023; Anil et al., 2023b; Touvron et al., 2023; Anthropic, 2024; Kaddour et al., 2023; Dubey et al., 2024) marked a significant advancement in generative models, enabling interaction with computing devices through natural language. This advancement rendered many earlier benchmarks and leaderboards obsolete (Laskar et al., 2023; Shen et al., 2023), leading to the compilation of more challenging and comprehensive tests. Among these benchmarks, Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021) has gained significant popularity: It assesses both the breadth and depth of language understanding capabilities of current LLMs across a diverse range of subjects, including mathematics, history, computer science, logic, law, etc.
Figure 1: Examples of erroneous instances from MMLU Virology, College Chemistry, and Human Aging.
However, the reliability of benchmarking results is only as robust as the quality of the dataset used. We find that, despite its popularity, MMLU suffers from numerous errors that can mislead evaluation and model comparison (L and Stapleton, 2023; Erenrich, 2023). These errors, which range from simple parsing and scraping mistakes to more complex issues related to context, interpretation, and dataset quality, compromise the reliability of MMLU as a benchmark. For example, we find that 57% of the analysed instances in the Virology subset contain errors, including the suggestion to send the American army to West Africa to prevent outbreaks of Ebola (see Fig. 1).
Therefore, in this study, we manually analyse the MMLU dataset using a novel error annotation protocol to construct MMLU-Redux: 14 human experts manually assessed and re-annotated 5,700 questions across all subsets of MMLU. We estimate that 6.49% of the questions are erroneous.
After our manual re-annotation effort, we study how the errors in MMLU impact LLM evaluation. First, we re-evaluated leading LLMs on MMLU-Redux, and found the performance metrics notably altered, changing their ranking. Furthermore, we both quantitatively and qualitatively analysed the errors to help understand how these errors impact LLM evaluation. MMLU-Redux can also be used as a strong benchmark for automatic error detection in NLP datasets, which would help scale up the review of benchmark datasets. Therefore, we also study whether LLMs can help with error detection, using prompting techniques (i.e., In-Context Learning (Brown et al., 2020), Chain of Thoughts (CoT) (Wei et al., 2022)), Retrieval Augmented Generation (RAG) (Lewis et al., 2020), and fine-tuning. We believe MMLU-Redux underscores the need for closely studying and reassessing the benchmarks used for evaluating NLP models.
@@ 2 What Is Wrong with MMLU?
The MMLU dataset has become a popular choice for evaluating the performance of NLP systems, owing to its extensive coverage of subjects collected from freely available online sources with the help of graduate and undergraduate students (Hendrycks et al., 2021). Despite the manual effort, MMLU still contains errors that are difficult to trace due to its under-documented annotation procedure. These errors have not yet been systematically studied, even though they were recently highlighted (L and Stapleton, 2023; Erenrich, 2023).
We identify various errors in MMLU questions, from simple parsing mistakes (e.g., the source answer is B, but MMLU labels it as C) to more complex issues, such as missing context. These errors occur randomly, and without detailed documentation, tracing their root causes is challenging.
This randomness and lack of traceability highlight the need for a standardised error categorisation to improve the reliability and accuracy of the MMLU dataset. Our approach involved developing a hierarchical annotation protocol of errors, which we used to develop MMLU-Redux: A manual annotation of all 57 subsets of MMLU, each containing 100 randomly selected samples (Section 3.1).
2.1 Error Detection Annotation Protocol
Figure 2: The error annotation protocol used for annotating instances in MMLU — annotators first assess whether the instance is unambiguous (Question Assessment, then whether there is a single valid answer, and then whether it matches with the ground truth answer from the dataset.
We develop a hierarchical annotation protocol to classify the various errors identified in MMLU into specific error types. Figure 2 illustrates our annotation protocol for categorising MMLU errors, while Figure 4 provides examples of each error category. We categorise errors into two primary groups: samples with errors in the clarity of the questions (Type 1, Question Assessment) and samples with errors in the ground truth answer (Type 2, Ground Truth Verification).
@@ 2.1 Error Detection Annotation Protocol
(2a) No Correct Answer: None of the options correctly answer the question. This error may arise when the ground-truth options are omitted to reduce the number of options from five to four.
@@
To create MMLU-Redux, we followed the proposed annotation protocol. We aim to ensure comprehensive coverage of the different error types for further experiments.
2.2 Heterogeneity of Errors in MMLU
Our annotation process reveals substantial variation in error types across subsets. Some subsets are affected by ambiguous questions, while others by incorrect ground truth labels. These differences may impact how MMLU results are interpreted and addressed. Notable irregularities in certain subsets include (a full list of subsets and their corresponding errors can be found in Appendix I):
@@ 2.2 Heterogeneity of Errors in MMLU
– Incorrect ground truth labels are prevalent within the Virology subset. Many of the incorrect labels are for relatively simple questions, such as identifying the description of a pandemic; this suggests errors may stem from problems parsing the original datasets (in most cases, the Human Virology textbook).
@@
– The questions were sourced from textbooks and standardised college-level exams. We identified erroneous questions resulting from simple parsing errors and unknown causes. For example, questions spanning multiple lines in the original source were often parsed incorrectly, leading to a part of the question being presented as the first option (Option A) and the exclusion of Option D. Furthermore, there were questions with ground truth labels that did not match the answers provided in the source without known cause.
@@
– The benchmark does not clearly distinguish between different jurisdictions, despite focusing on U.S. law.
@@
– The dataset contains many questions with incorrect answers. These are mostly sourced from the ‘Oxford Learning Link’ website. It is unknown what causes the inaccuracies. For example, one question states that $(F\wedge L)\wedge\neg C$ is correct, but $F\wedge L\wedge\neg C$ is not, even though these two formulas are clearly equivalent.
@@
These irregularities highlight error patterns in MMLU. We want to highlight one type of error that is especially challenging to catch, namely unspecified context that is needed to properly answer the question. For example, the Professional Law and Accounting datasets assume US jurisdiction and practices which are not specified in the question and may become outdated if standards change. Many subjects also display a US- and Western-centric bias: the Virology dataset refers to “the Latino population” in a US context, and the Human Aging dataset mentions an unspecified survey of older adults.
Figure 3: Statistics of the error types detected in 100 randomly sampled and manually annotated data from the 30 most erroneous MMLU subjects. In the Virology subject, we found that 57% of the analysed instances contain some forms of errors, such as wrong ground truth (33%), multiple correct answers (4%), or unclear questions (14%). For detailed numbers of all subsets, see Appendix F.
Figure 4: An example for every leaf node in the error annotation protocol shown in Fig. 2 from the MMLU dataset.
@@ 3 MMLU-Redux: A Correct MMLU Subset
We create MMLU-Redux by manually labelling a subset of MMLU questions with their corresponding error types. To this end, we follow the annotation protocol introduced in Section 2.1. For more accurate annotations, we confirmed the error detection by finding the samples’ original source wherever it was available. However, at present, the correct answers suggested by the annotators are not used to replace the existing MMLU labels.
In the following, we analyse the error statistics of MMLU-Redux and use MMLU-Redux to re-evaluate the performance of LLMs. In Section 4, we explore the possibility of using MMLU-Redux to improve the overall quality of MMLU by automatically fixing the identified errors.
@@ 3.1 Analysis of MMLU-Redux
We present the percentage of error types in Fig. 3, with detailed numbers available in Appendix F. In our analysis, we find that more than 9% of the examples are incorrect, suggesting a substantial presence of errors in the MMLU. Especially, we find that more than 57% examples in Virology contain errors, where 30% examples have a wrong ground truth, and 15% are unclear questions. Moreover, we also observe significant error percentages in other disciplines: more than 20% examples in Logical Fallacies and College Chemistry are wrong, and more than 10% examples in Professional Law, Business Ethics, Formal Logic, Human Aging, Global Facts, Machine Learning, Miscellaneous and Public Relations are wrong. Such error proportions could lead to inaccurate comparisons and invalid rankings of LLM models. Using stratified sampling, we estimate that 6.49% of the questions in the complete MMLU dataset contain errors.
Figure 5: Exact Match (EM) for each model on all subjects combined (top left), and the seven subjects with the most errors in our MMLU-Redux data. Original EM represents performance measured across all instances regardless of error presence, while Correct and Erroneous EM reflect results on correct instances and error instances, respectively.
To better understand how these errors impact the performance of models on the MMLU, we compare the performance between erroneous instances and correct instances across the seven subjects identified as having the most errors (Virology, Logical Fallacies, College Chemistry, Professional Law, Business Ethics, Formal Logic, and Human Aging) in Fig. 5.
Despite the general trend indicating a performance decline among erroneous instances, we also observed cases where performance was similar or even higher in erroneous instances (Professional Law and Formal Logic). Given that LLMs should be unable to yield answers that are correct with respect to the original MMLU ground truth on these erroneous instances, this may be evidence of memorisation, suggesting that these MMLU instances were learned during the pretraining.
Table 1: Inter-annotator agreement (Cohen’s Kappa) analyses for the five erroneous subjects of MMLU-Redux. “All Error Type” is calculated based on whether the specific error types match, while “Binary Error Type” is calculated based on whether an error is present or not.
Subject
Cohen’s Kappa
(All Error Type)
Cohen’s Kappa
(Binary Error Type)
Virology 0.67 0.67
Logical Fallacies 0.73 0.71
College Chemistry 0.92 0.95
Formal Logic 0.96 1.0
Human Sexuality 0.64 0.64
We also assessed the reliability of the manual annotations in MMLU-Redux by calculating Cohen’s Kappa (Cohen, 1960) among three annotators for the five erroneous subjects (virology, logical fallacies, college chemistry, formal logic, and human sexuality). The results shown in Table 1 demonstrate a generally high level of agreement across most subjects, with Cohen’s Kappa values exceeding 0.6 for both “All Error Type” (whether the specific error types match) and “Binary Error Type” (whether an error is present or not). Additionally, the minimal difference in scores between “All Error Type” and “Binary Error Type” indicates that our error annotation protocol used for annotating (Section 2.1) is highly effective.
@@ 3.2 Re-Evaluating the State-of-the-Art LLMs
To assess how the corrected dataset impacts the performance of existing state-of-the-art LLMs, we re-evaluate them on the five subjects with the highest number of errors (Virology, Logical Fallacies, College Chemistry, Professional Law, and Human Sexuality) in MMLU-Redux.
@@
In Table 2, we compared the performance of models when using all instances of MMLU-Redux to the performance when using only correct instances without errors to see if there are any changes in the rankings due to this. The results clearly demonstrate that, at least for subjects with a high number of detected errors, these issues are significant enough to affect the results, leading to changes in model rankings. For example, in the Virology subset, Llama 3.1 Instruct Turbo (405B) ranked 16th when considering all Virology instances, but ranked first when only correct instances were used. On the other hand, in the Human Sexuality subset, GPT-4 (0613) achieved an exact match score of 0.91 across all instances (ranking 5th). However, when considering only the correct instances, the exact match score drops significantly to 0.43, placing it last among the top 10 models. Since MMLU is an important benchmark for evaluating model performance, this indicates that errors in MMLU are a critical issue.
@@
Table 2: Comparison of model performance and ranking changes when using overall instances (as per HELM Liang et al. (2023)) versus correct instances of MMLU-Redux on five most erroneous subjects. Numbers in parentheses denote ranks, teal an improvement in rank, red a deterioration in rank, and black an unchanged rank.
@@
Model Virology Logical Fallacies College Chemistry Professional Law Human Sexuality
Claude 3.5 Sonnet (20240620) 0.60 (1) $\rightarrow$ 0.91 (5) 0.93 (1) $\rightarrow$ 0.96 (5) 0.59 (9) $\rightarrow$ 0.73 (4) 0.75 (1) $\rightarrow$ 0.77 (1) 0.94 (1) $\rightarrow$ 0.98 (1)
Claude 3 Opus (20240229) 0.58 (12) $\rightarrow$ 0.88 (8) 0.90 (4) $\rightarrow$ 0.96 (5) 0.60 (5) $\rightarrow$ 0.72 (5) 0.72 (4) $\rightarrow$ 0.72 (3) 0.91 (5) $\rightarrow$ 0.96 (2)
Llama 3.1 Instruct Turbo (405B) 0.57 (16) $\rightarrow$ 0.93 (1) 0.92 (2) $\rightarrow$ 0.96 (5) 0.60 (5) $\rightarrow$ 0.76 (1) 0.70 (6) $\rightarrow$ 0.72 (3) 0.86 (20) $\rightarrow$ 0.91 (9)
GPT-4o (2024-05-13) 0.60 (3) $\rightarrow$ 0.91 (5) 0.88 (6) $\rightarrow$ 0.99 (2) 0.61 (4) $\rightarrow$ 0.71 (7) 0.72 (3) $\rightarrow$ 0.70 (5) 0.91 (5) $\rightarrow$ 0.96 (2)
Gemini 1.5 Pro (001) 0.55 (28) $\rightarrow$ 0.91 (5) 0.90 (4) $\rightarrow$ 0.99 (2) 0.62 (2) $\rightarrow$ 0.72 (5) 0.67 (9) $\rightarrow$ 0.67 (7) 0.37 (55) $\rightarrow$ 0.94 (6)
GPT-4 (0613) 0.60 (3) $\rightarrow$ 0.86 (10) 0.87 (11) $\rightarrow$ 0.99 (2) 0.55 (18) $\rightarrow$ 0.75 (3) 0.73 (2) $\rightarrow$ 0.68 (6) 0.91 (5) $\rightarrow$ 0.43 (10)
Qwen2 Instruct (72B) 0.56 (24) $\rightarrow$ 0.88 (8) 0.91 (3) $\rightarrow$ 0.96 (5) 0.65 (1) $\rightarrow$ 0.68 (8) 0.66 (10) $\rightarrow$ 0.74 (2) 0.89 (11) $\rightarrow$ 0.94 (6)
GPT-4 Turbo (2024-04-09) 0.60 (1) $\rightarrow$ 0.93 (1) 0.87 (11) $\rightarrow$ 1.00 (1) 0.53 (22) $\rightarrow$ 0.76 (1) 0.67 (8) $\rightarrow$ 0.63 (9) 0.90 (9) $\rightarrow$ 0.93 (8)
Gemini 1.5 Pro (0409 preview) 0.58 (10) $\rightarrow$ 0.93 (1) 0.86 (18) $\rightarrow$ 0.92 (10) 0.58 (13) $\rightarrow$ 0.67 (9) 0.64 (13) $\rightarrow$ 0.61 (10) 0.40 (56) $\rightarrow$ 0.95 (5)
Llama 3.1 Instruct Turbo (70B) 0.58 (12) $\rightarrow$ 0.93 (1) 0.84 (27) $\rightarrow$ 0.96 (5) 0.59 (9) $\rightarrow$ 0.64 (10) 0.67 (7) $\rightarrow$ 0.65 (8) 0.86 (20) $\rightarrow$ 0.96 (2)
3.3 Correlation between Exact Match Score and Data Quality
Figure 6 shows the correlation between the performance of the ten best-performing LLMs on MMLU (as ranked by HELM Liang et al. (2023)) and the ratio of non-erroneous instances on MMLU-Redux. Each point in the scatter plot represents a subject in MMLU. As expected, LLMs perform better in subjects with fewer errors, supporting the link between data reliability and model performance. Ideally, LLMs excel with non-erroneous instances and struggle with erroneous ones.
Outliers, however, reveal deviations from this trend, where models perform better or worse than expected for certain subjects. These deviations suggest other factors influence performance, requiring further investigation. For example, some LLMs perform well despite high error rates (in the Logical Fallacies subject), hinting at potential subject-specific advantages. We further explore these anomalies and their implications for model robustness in Section 3.1.
Figure 6: Correlation between the performance of 10 best-performing LLMs on MMLU (as per HELM Liang et al. (2023)) and the ratio of non-erroneous instances on MMLU-Redux. Each dot represents one MMLU subject.